he Foundational 15 [LO2-1, LO2-2, LO2-3, LO2-4]

[The following information applies to the questions displayed below.]

Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments—Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 10,000 $ 15,000 $ 25,000
Estimated variable manufacturing overhead per machine-hour $ 1.40 $ 2.20
Job P Job Q
Direct materials $ 13,000 $ 8,000
Direct labor cost $ 21,000 $ 7,500
Actual machine-hours used:
Molding 1,700 800
Fabrication 600 900
Total 2,300 1,700

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Sweeten
Answer 1
Plant wide predetermined OH Rate Molding Fabrication Total Note
Machine Hours        2,500.00        1,500.00     4,000.00 A
Variable OH Rate                1.40                2.20 B
Variable OH cost        3,500.00        3,300.00     6,800.00 C=A*B
Fixed Manufacturing overhead      10,000.00      15,000.00 25,000.00 D
Total OH 31,800.00 E=C+D
Plant wide predetermined OH Rate             7.95 F=E/A
Answer 2 Job P Job Q Note
Actual Machine hours        2,300.00        1,700.00 G
Plant wide predetermined OH Rate                7.95                7.95 See F
Overhead applied     18,285.00      13,515.00 H=G*F
Answer 3 Amount $ Note
Direct Materials      13,000.00
Direct Labor      21,000.00
Overhead applied      18,285.00 See H
Manufacturing cost assigned to Job P     52,285.00 I
Answer 4 Amount $
Manufacturing cost assigned to Job P      52,285.00 See I
Number of units             20.00 J
Unit product cost        2,614.25 K=I/J
Answer 5 Amount $
Direct Materials        8,000.00
Direct Labor        7,500.00
Overhead applied      13,515.00 See H
Manufacturing cost assigned to Job Q     29,015.00 L
Answer 6 Amount $
Manufacturing cost assigned to Job Q      29,015.00 See L
Number of units             30.00 M
Unit product cost           967.17 N=L/M
Answer 7 Job P Job Q Note
Manufacturing cost assigned      52,285.00      29,015.00 O=K,N
Mark up at 80%      41,828.00      23,212.00 P=O*80%
Total price for the job     94,113.00      52,227.00 Q=P+O
Number of units             20.00              30.00 R=J,M
Selling price per unit        4,705.65        1,740.90 S=Q/R
Answer 8 Amount $ Note
Manufacturing cost assigned to Job P      52,285.00 See K
Manufacturing cost assigned to Job Q      29,015.00 See N
Cost of goods sold     81,300.00
Answer 9
Calculation of predetermined OH Rate Molding Fabrication Note
Machine Hours        2,500.00        1,500.00 See A
Fixed Manufacturing overhead      10,000.00      15,000.00 See D
Fixed Manufacturing overhead rate                4.00              10.00 T=D/A
Variable OH Rate                1.40                2.20 See B
Departmental predetermined OH Rate                5.40              12.20 U=T+B
Answer 10- Job P Molding Fabrication Total Note
Machine hours used        1,700.00            600.00     2,300.00 V
Departmental predetermined OH Rate                5.40              12.20 See U
Overhead applied in Job P        9,180.00        7,320.00 16,500.00 W=V*U
Answer 10- Job Q Molding Fabrication Total
Machine hours used           800.00            900.00     1,700.00 X
Departmental predetermined OH Rate                5.40              12.20 See U
Overhead applied in Job Q        4,320.00      10,980.00 15,300.00 Y=X*U